# Robustness-aware Automatic Prompt Optimization

### v2

@@ Abstract
The performance of Large Language Models (LLMs) depends on the quality of prompts and the semantic and structural integrity of the input data. However, existing prompt generation methods primarily focus on well-structured input data, often neglecting the impact of perturbed inputs on prompt effectiveness. To address this limitation, we propose BATprompt (By Adversarial Training prompt), a novel method for prompt generation designed to withstand input perturbations (such as typos in the input). Inspired by adversarial training techniques, BATprompt demonstrates strong performance on a variety of perturbed tasks through a two-step process: adversarial perturbation and iterative optimization on unperturbed input via LLM. Unlike conventional adversarial attack methods, BATprompt does not need access to model parameters and gradients. Instead, BATprompt leverages the advanced reasoning, language understanding and self reflection capabilities of LLMs to simulate gradients, guiding the generation of adversarial perturbations and optimizing prompt performance. We evaluate BATprompt on multiple datasets across both language understanding and generation tasks. The results indicate that BATprompt outperforms existing prompt generation methods, delivering superior robustness and performance under diverse perturbation scenarios. The code is available at: https://github.com/vanpe20/BATprompt.
Robustness-aware Automatic Prompt Optimization
Zeru Shi1, Zhenting Wang2, Yongye Su3, Weidi Luo4, Hang Gao2, Fan Yang5, Ruixiang Tang2, Yongfeng Zhang2, 1 Dalian University of Technology, 2 Rutgers University, 3 Purdue University, 4 The Ohio State University, 5 Wake Forest University
@@ 1 Introduction
LLMs can perform a wide range of tasks including text classification, summarization, generation Abburi et al. (2023); Ge et al. (2024); Laban et al. (2023). With their broad applicability, researchers have begun exploring strategies to enhance the performance of LLMs on these tasks by effectively activating their capabilities.
@@
Figure 1: An example of adding a perturbation to the input and its result in language understanding task, where the top one indicates that the prompt gets the correct output under normal input, the middle one indicates that the prompt gets the wrong result under perturbed input, the bottom indicates the BATprompt gets correct result under perturbed input.
To address this issue, researchers have begun designing more effective prompts to enhance the performance of LLMs. To reduce the burden of manual prompt design, LLMs are now widely utilized to optimize prompts and generate improved candidates. Some existing methods include fine-tuning the performance of LLMs (Ouyang et al., 2022; Ziegler et al., 2019), optimizing the prompt by generating CoT (Chain-of Thought) (Shum et al., 2023; Zhang et al., 2022), and optimizing the prompt by guiding the inference ability of LLMs (Wang et al., 2023b; Liu et al., 2024). However, despite the success of these methods, a critical issue is overlooked: in real-world scenarios, task inputs frequently contain errors, such as typos, vague expressions, or inaccuracies. Under such conditions, the prompts generated by current prompt optimization techniques may not exhibit sufficient robustness to handle such imperfections in the input data. For example, in  Figure 1, the normal prompt will make LLMs get wrong answer in the classification task when input text has typos.
To enhance prompt robustness for task performance, many researchers have turned to data augmentation, adding perturbed text to the training data and training the model on this augmented dataset to produce robust prompts. However, our experimental results indicate that these prompt generation approaches frequently yield prompts that lack robustness against input perturbations (see §5.6). In some tasks, the performance of these prompts even declines, likely due to the excessive diversity of perturbation types introduced during augmentation. We contend that while basic data augmentation methods can be beneficial in certain contexts, they may inadvertently introduce noise that interferes with effective prompt generation and limits generalization to unseen perturbations.
To enable the model to better identify disturbances in the data and enhance the generalization ability of the generated prompts, We introduce a novel prompt optimization method, called BATprompt, designed to generate discrete instructions resilient to perturbations. Our approach takes inspiration from adversarial training Madry (2017), which is widely recognized for improving neural network robustness by exposing the model to adversarial attacks during training. We found that the concept of adversarial attacks can also be applied to prompt optimization, making the generated prompts more resilient to perturbed inputs. In the attack phase of our approach, we use gradient updates to deliberately induce adversarial inputs that degrade LLM performance. During the adversarial training stage, we implement two distinct optimization modes tailored to address different types of attacks, iteratively refining the prompts using gradient-based guidance. This ensures that the resulting prompts maintain robustness under diverse conditions. Our main contributions are as follows:
  • We propose an adversarial attack-inspired method to perturb the input text without using model gradient, aiming to identify the vulnerabilities of existing prompt optimization techniques to such text perturbations.
  • We introduce a novel framework BATprompt for generating prompts by adversarial prompt optimization, which are resilient to input perturbations. Using an adversarial training framework, we harness gradient-based techniques to drive both the attack and prompt generation processes. The resulting prompts show robust performance in many perturbation types.
  • We present a perturbation dataset, which includes various tasks such as language understanding and language generation. For each task, we added different levels of perturbation, such as character level, word level and sentence levels.
  • Experimental results demonstrate that the proposed method achieves exceptional performance across multiple tasks under various types of perturbations, validating the robustness of BATprompt in addressing these challenges.
2 Relarted Work
2.1 Adversarial Training
Adversarial Attack: Numerous adversarial attack methods (Goodfellow et al., 2014; Madry, 2017; Carlini and Wagner, 2017), have been developed for computer vision and traditional NLP tasks. Among these, FGSM Goodfellow et al. (2014) and PGD Madry (2017) are utilized to add adversarial attack on data. Nowadays, people begin to study adversarial attacks against LLMs, which involve adding slight modifications to natural language input, such as spelling errors, synonym substitutions, character substitutions, or out-of-order, to trick LLMS into making wrong predictions or generation. Zhou et al. (2024c) add attacks on math solving problems in LLMs. Zhu et al. (2023) classifies the hint attacks on LLM into four categories. Specific examples of different classes of attacks are given in Xu et al. (2024). Among them, word-level Wang et al. (2023a); Zhou et al. (2024b) and sentence-level Gu et al. (2023); Dong et al. (2021) attacks are more common. Zhou et al. (2024b) proposed a word-level attack based on the classification task without changing the semantics.
Adversarial Training: FGSM Goodfellow et al. (2014) proposed the concept of adversarial training, which optimizes the model after adversarial attacks on the data. Miao et al. (2024) proposes a method for an effective adversarial attack on T2M. In LLMs,  Raina et al. (2024) presents the unrobustness of Judge-LLM against adversarial attacks, which leads to the inflation of LLM scores. Yao et al. (2023a); Kumar et al. (2023) proposed adversarial attacks on different aspects of LLM and proposed several defense methods, and Yao et al. (2023a) proposed several methods to defend against malicious attacks in prompts. Kumar et al. (2023) proposes an adversarial hallucination attack and proposes a defense strategy Sheshadri et al. (2024); Xhonneux et al. (2024); Lin and Zhao (2024). The adversarial training method is used to improve the robustness of LLM. Sheshadri et al. (2024) proposed a method ReFAT to simulate input attacks and defend against them by refusal feature ablation. Xhonneux et al. (2024) Latant Adversarial Traning is used to improve the robustness of LLM against Jalibreak. Lin and Zhao (2024) proposed a defense technique called LLAMOS to enhance the adversarial robustness of LLM through adversarial attacks of text perturbation.
2.2 Prompt Optimization
The manually designed LLM prompt method sometimes does not make LLM perform better in performing tasks, so the prompt optimization method arises. Li et al. (2023) proposes a fine-tuning method based on context ordering and probability ordering. Beyond fine tuning, reinforcement learning is also a great optimization  Ma et al. (2023); Sun et al. (2023); Yao et al. (2023b), and  Ma et al. (2023) proposes Eureka, which generates a reward function that outperforms human experts. Sun et al. (2023) optimizes the arithmetic reasoning ability of large models by Prompt-OIRL. RLprompt Deng et al. (2022) introduces a reword mechanism to generate better prompts. At present, LLM optimization has become the main method of prompt optimization Shum et al. (2023); Zhang et al. (2022); Yang et al. (2023)Zhou et al. (2022) proposed APE to use LLM to automatically generate prompts according to input and output. Pryzant et al. (2023) proposed to use LLM self-correction as pseudo-gradient to optimize prompts. Guo et al. (2023) uses genetic algorithm to optimize LLM prompts through heredity and mutation. Jin et al. (2024) search the impact of step length in activating LLMs.  Zhou et al. (2024a) use adversarial attack to generate more success jailbreak to deceive LLMs and add some suffix after the input text to resist jailbreak. Additionally, most existing methods generate optimized prompts using clean datasets, overlooking the inevitable perturbations present in natural language inputs. These perturbations can negatively affect the effectiveness of the prompts. Our method addresses this limitation by proposing a strategy that emphasizes maintaining robustness against such perturbed inputs.
@@ 2.2 Prompt Optimization
Figure 2: Heat maps showing the magnitude of the impact of each perturbation on different types of tasks, where darker colors indicate a stronger impact of the perturbation on this type of task, and vice versa a lower one. No color indicates no effect or a positive effect
Figure 3: The workflow of an iteration of BATprompt. The Adversarial Attack Phase is used to generate the adversarial samples. The Adversarial Optimization Phase is used to generate optimized prompt. rk means score ranking. For example, rk1 means that the score is ranked first in all outputs.
3 Perturbation Types and Task Vulnerability
@@
We adopt a widely used approach following Xu et al. (2023), which form the basis for the perturbations in our experiments (See Table 1). These disturbances are categorized into two groups: P1 and P2. P1 includes: C1, C2, C3, and S1. The perturbations in this category primarily introduce typographical errors and non-sensical strings into the text, without significantly altering the underlying semantic structure of the sentence. P2 includes: W1, W2, W3, S2, and S3. Unlike P1, these perturbations alter the semantic structure of sentences by modifying their syntactic composition, while maintaining their core meaning. These modifications can result in a bias in how LLMs interpret the input, potentially affecting their understanding.
To assess the weakness of different language tasks on these nine types of perturbations, we added each perturbation to the dataset of different tasks and evaluated the performance of Manual Instructions Bach et al. (2022) and Natural Instructions Mishra et al. (2021) on these datasets. The results of the impact of the perturbation on the task are shown in  Figure 2, where darker colors indicate greater impact and vice versa. For each task, we select perturbations that have a significant impact as potential candidates and disregard those with a positive or negligible effect. It can be concluded that perturbations of class C3 have no negative impact on any task, whereas perturbations of class P2 consistently affect all tasks to varying degrees. We observe that character-level perturbations in Simplification tasks can positively impact task performance. We believe this is because typo perturbations encourage LLMs to focus more on the core meaning of the input, aligning better with the objective of the simplify task. The specific experimental results and analysis, and the details of the experimental settings are in Appendix A.
4 Methodology
To enhance the robustness of auto-prompt under perturbation, we propose introducing adversarial training into this process. Traditional adversarial training depends on gradients to generate adversarial examples and guide training, but most LLMs operate via black-box APIs where gradients and parameters are inaccessible. Thus, we introduce BATprompt, that employs guided information instead of gradients. By leveraging this guidance, LLMs generate new prompts through adversarial attacks and iterative optimizations, enhancing their robustness in context learning.
Algorithm 1 The algorithm flow of BATprompt
1:Initial prompts $p$, m of adversarial attack gradients $g_{adv}$ and optimization gradient $g_{opt}$, adversarial attack operation $f_{adv}(\cdot)$, optimization operation $f_{opt}(\cdot)$. adversarial attack constraints $\mathcal{D}(\cdot)$. Sorting function $\mathcal{S}(\cdot)$.
2:for $t=1$ to $T$ do
3:     Random Select: Num of texts $T^{i}$ (i $\in$ N) are randomly selected from the unperturbed dataset.
4:     for $j=1$ to num do
5:         Adversarial Attack: $T^{{}^{\prime}}\leftarrow f_{adv}(T^{i},g_{adv}^{k})$, where i $\in$ (1, N), k $\in$ (1, m).
6:         Select Worst Text: $\mathcal{D}(T^{{}^{\prime}},T^{i})<\epsilon$, $Min$ $\mathcal{S}(T^{{}^{\prime}})$.
7:     end for
8:     Generate Gradient: Generate optimized gradient $g_{opt}$ through the perturbed text.
9:     Optimize Prompt: $p\leftarrow f_{opt}(p,g_{opt})$
10:     Select Best Prompt: $Max$ $\mathcal{S}(p)$
11:end for
12:Return: Returns the best prompt $p_{best}$ with the highest rating on the perturbed task.
4.1 Threat Model
Adversarial Perturbation Goal. Like traditional adversarial attacks, this phase aims to induce incorrect outputs from model $\mathcal{M}(\cdot)$ by adding a perturbation to the input. Specifically, given a clean sample $x$ and a perturbed input $x^{\prime}$, we aim for$\mathcal{M}(x^{\prime})=y$ $\mathcal{M}(x^{\prime})=y^{\prime}$, where $y$ and $y^{\prime}$ represent the corresponding outputs, and $y^{\prime}\neq y$. The perturbation must meet the following two conditions:
  • Maintain the semantic similarity and the structural similarity to the original text.
  • Degrade the performance of the large language model (LLM) on the task, such that the model’s performance on the perturbed input is worse than on the clean input.
Prompt Optimization Goal. Users typically have typos or semantic ambiguities when typing, especially when dealing with large volumes of perturbed text. Therefore, the goal of prompt optimization is to train robust instructions that can still ensure good task performance, even when confronted with perturbed text inputs.
4.2 Framework of BATprompt
BATprompt follows a similar principle to existing gradient-based adversarial attack methods. It first iteratively adds attacks within the specified range to undisturbed inputs by adjusting them along the gradient direction to maximize disruption. The adversarial samples generated are then used for training based on a defined loss function. The workflow of our method is shown in Figure 3. The details of the specific implementation in LLM are shown in Figure 11. BATprompt contains many iterations and every iterations include two key components:
  • Adversarial Attack Phase. BATprompt begins with manually crafted prompts and unperturbed text as the initial input. In  section 3, we categorize nine distinct perturbations into two types and design tailored adversarial attack methods for each category. These methods generate adversarial samples that simulate various perturbation scenarios, allowing the subsequent optimization algorithms to build robustness across all types of perturbations. This approach ensures that the model remains effective when faced with diverse perturbations.
  • Adversarial Optimization Phase. At this stage, we employ an iterative optimization method that incorporates a "gradient" mechanism to guide the LLM in refining the prompts. For each generated prompts, we select the candidate with the best performance on the validation set and retain it for the next iteration.
4.3 Adversarial Attack Phase
@@
At this stage, we employ an adversarial attack method inspired by FGSM Goodfellow et al. (2014), which uses a fixed gradient direction to enable the model to rapidly generate adversarial samples. In BATprompt, we introduce various types of perturbation-specific guide words as a fixed gradient $g$, and apply perturbations to $n$ samples $x$, randomly selected from the original unperturbed dataset. $y$ is the standard output. The performance of the original prompt on the perturbed inputs is evaluated using a loss function $\mathcal{L}_{adv}$. The score is minimized when perturbed. After generating the perturbed sample prompt $p$, we compute either its Levenshtein distance or semantic similarity $||x^{\prime}-x||<\epsilon$, depending on the nature of the perturbation. The perturbations are defined as follows:

### v3

# Auto-Prompt Generation is Not Robust: Prompt Optimization Driven by Pseudo Gradient

@@ Abstract
While automatic prompt generation methods have recently received significant attention, their robustness remains poorly understood. In this paper, we introduce PertBench, a comprehensive benchmark dataset that includes a wide range of input perturbations, designed to systematically evaluate the robustness of current auto-prompting techniques. Our analysis reveals substantial vulnerabilities in existing prompt generation strategies, where even minor modifications to the prompt can lead to significant differences in model output. To address this issue, we propose PGO, a gradient-free prompt generation framework that leverages perturbation types as pseudo-gradient signals to guide LLMs in producing more robust prompts. In contrast to existing methods that assess prompt quality only on clean, well-structured inputs, our approach explicitly emphasizes robustness under noisy and perturbed conditions. Extensive experiments across diverse tasks and multiple LLMs show PGO consistently outperforms previous methods in maintaining performance under input perturbations.
Auto-Prompt Generation is Not Robust:
Prompt Optimization Driven by Pseudo Gradient
Zeru Shi1, Zhenting Wang1, Yongye Su2, Weidi Luo3, Hang Gao1, Fan Yang4, Ruixiang Tang1, Yongfeng Zhang1, 1 Rutgers University, 2 Purdue University, 3 Georgia University, 4 Wake Forest University
@@ 1 Introduction
Figure 1: An example of adding a perturbation to the input and its result in classification task, where the top one indicates that the prompt gets the correct output under normal input, the middle one indicates that the prompt gets the wrong result under perturbed input, the bottom indicates the PGO gets correct result under perturbed input.
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide spectrum of tasks. Abburi et al. (2023); Ge et al. (2024); Laban et al. (2023). Given their broad applicability, a growing body of research has focused on strategies to more effectively elicit and amplify these capabilities. Among them, prompt engineering has emerged as a central technique, with recent advances increasingly relying on LLMs themselves to automate the prompt optimization process. Representative approaches include instruction fine-tuning Ouyang et al. (2022); Ziegler et al. (2019), Chain-of-Thought (CoT) prompting Shum et al. (2023); Zhang et al. (2022), and inference-guided prompt generation Wang et al. (2023b); Liu et al. (2024). Despite their success on clean, well-structured inputs, these methods share a key limitation: they assume idealized input conditions, and thus fail to account for the noisy, imperfect inputs commonly encountered in real-world scenarios, such as typographical errors, ambiguous phrasing, or factual imprecision. These input perturbations can significantly degrade model performance, revealing a critical robustness gap in current prompt optimization techniques. As shown in Figure. 1, even subtle input corruptions (e.g., a minor typo) can mislead an LLMs and trigger incorrect predictions in tasks like classification.
Figure 2: Comparison of the execution results of all different methods on PertBench for the three tasks.
@@
In this work, we introduce PertBench, a comprehensive benchmark designed to evaluate the robustness of prompt-based methods under diverse input perturbations. Our benchmark spans three representative natural language processing tasks, text simplification, summarization, and classification—capturing a wide spectrum of practical use cases. In contrast to prior datasets such as NoiseBench Merdjanovska et al. (2024); Moradi and Samwald (2021); Dong et al. (2023), which primarily focus on character-level noise within a single task or short form content (e.g., dialogues), PertBench provides broader coverage across multiple granularity levels, including character-level, word-level, and sentence-level perturbations. Moreover, we extend the perturbation setting to long form text and ensure a uniform and comprehensive perturbation strategy across all tasks. The specific categories of perturbations are summarized in Table 1. This design enables a more systematic and rigorous evaluation of prompt robustness across tasks and text lengths. Our results demonstrate that existing prompt generation methods experience substantial performance degradation under perturbed inputs, exposing their vulnerability and limited generalization.
To mitigate this issue, we propose PGO (Pseudo Gradient Optimization), a gradient-free framework for robust discrete prompt optimization. Unlike conventional approaches that rely on model gradients to update prompts, PGO leverages perturbation information as guidance to guide prompt refinement. We categorize perturbations into two types P1 and P2 depending on whether they cause significant semantic shifts. Correspondingly, PGO introduces two complementary optimization strategies tailored to these categories and iteratively refines prompts via feedback-driven updates. Through this design, PGO effectively enhances prompt robustness against a broad range of input disturbances. Our main contributions are summarized as follows:
  • We present PertBench, a comprehensive benchmark for evaluating prompt robustness under diverse input perturbations. It includes nine perturbation types across three core task categories for systematic robustness assessment.
  • We propose PGO, a pseudo gradient–based prompt optimization framework that enhances robustness across perturbations without requiring model internals or true gradients.
  • Extensive experiments show that existing methods degrade notably on PertBench, while PGO consistently improves robustness across tasks and perturbation settings.
2 Benchmark Construction
2.1 Details of Benchmark
2.1.1 Basic Datasets
In the classification tasks, we focus on six datasets to apply and test perturbations. These include sentiment classification datasets: SST-2 Socher et al. (2013), CR Hu and Liu (2004), SST-5 Socher et al. (2013), and MR PaNgB (2005), as well as topic classification datasets: AG’s News Zhang et al. (2015) and TREC Voorhees and Tice (2000). In language generation tasks, we utilize the Asset Alva-Manchego et al. (2020) for text simplification, which includes multiple benchmarks for reference translations. For the text summarization task, we use the XSum Narayan et al. (2018), which consists of concise summaries generated from longer texts. More details of PertBench in the Appendix A.
2.1.2 Perturbation Types
@@
We adopt a widely used approach following Xu et al. (2023), which form the basis for the perturbations in our experiments (See Table. 1). These disturbances are categorized into two groups: P1 and P2. P1 introduce typographical errors and non-sensical strings into the text, without significantly altering the underlying semantic structure of the sentence. P2 alter the semantic structure of sentences by modifying their syntactic composition, while maintaining their core meaning. These modifications can result in a bias in how LLMs interpret the input, potentially affecting their understanding. We compare the similarity between the perturbed data set and the original data set, and the experiment shows that we complete the perturbation of the data while maintaining the original similarity. The results in the Appendix A.
2.2 Benchmark Construction
@@
When constructing the dataset, we apply nine different perturbations to each type of task to construct separate perturbed datasets. To be more specific, we introduce various types of perturbation-specific guide words as a fixed gradient $g$, and use it to guide the perturbations on $x$. $y$ is the standard output. The performance of the original prompt on the perturbed inputs is evaluated using a loss function $\mathcal{L}_{adv}$. The score is minimized when perturbed. After generating the perturbed sample prompt $p$, we compute either its Levenshtein distance or semantic similarity $x^{\prime}-x||<\epsilon$, depending on the nature of the perturbation. The perturbations are defined as follows:
@@
For both types of perturbation, we adopt an iterative interference approach. At each step, we select the sentence that affects the outcome while exhibiting the highest semantic or structural similarity to the original sentence. Our optimization objective is formulated as follows:
$x^{\prime}=x_{0}+\mathcal{P}(x_{0},g)+\mathcal{P}(x_{1},g)+\ldots+\mathcal{P}(x_{n-1},g),$ (2)
Where $\mathcal{P(\cdot)}$ denotes that the given input $x$ is perturbed under the guide $g$, $g$ $\in$ {P1, P2}. $x_{n}$ denotes the text generated under each iteration. After each iteration, we compute the Levenshtein distance or semantic similarity between all perturbed texts and their corresponding source texts, selecting the resulting text as input for the next iteration. However, such iterative perturbation is generally only effective for short texts, such as classification or simplification tasks. When applied to long-text inputs, such as text summarization, this approach faces two main issues: The model exhibits weak sensitivity to P1-type perturbations. The perturbation model for P2-type perturbations tends to produce shorter outputs, resulting in deviations from the original intent. To address these issues, for long-text inputs, we perturb only one sentence at a time. This sentence is randomly selected from the text in each iteration like $s_{i}$ $\in$ $\{s_{1},s_{2},...,s_{n}\}$. $s$ denotes each sentence in the long text. This strategy preserves the accuracy of the model’s response when perturbations are introduced and helps maintain the overall semantic relevance of the generated text.
2.3 Results Analysis and Task Vulnerability
Figure 3: Heat maps showing the magnitude of the impact of each perturbation on different types of tasks, where darker colors indicate a stronger impact of the perturbation on this type of task, and vice versa a lower one. No color indicates no effect or a positive effect.
To evaluate the effectiveness of our constructed dataset in generating prompt perturbations, we tested the performance of all prompts used in the Baseline model under the PertBench framework. The detailed results are presented in the experimental section, alongside those of our proposed method for comparison. To provide a clearer understanding of model vulnerabilities, we visualize the performance of each baseline across all perturbations using radar plots in Figure. 2. These plots reflect the outcomes on three tasks: the first three graphs correspond to different metrics for the text summarization task; the fourth graph presents results for the text simplification task; and the fifth graph shows results for the text classification task. From the visualizations, it is evident that most perturbations degrade task performance, indicating their effectiveness in identifying vulnerabilities. However, not all perturbations have a negative impact-some even improve performance. For instance, in the summarization task, perturbations C3 and W2 do not reduce model performance. In the simplification task, perturbations C1, C2, and C3 show no negative effect. Similarly, for the classification task, C3 and S1 yield positive effects in many cases. To further illustrate these findings, we include heatmaps showing the effects of various perturbations on manual prompts in Figure 3. In these heatmaps, the numerical values represent the proportion of change compared to the original (unperturbed) input. The results clearly show that certain perturbations, particularly C3, have minimal or no negative impact across all three tasks. This suggests that the prompts involved in C3 are particularly robust to perturbations. We hypothesize that this is because the presence of input noise encourages the LLM to focus more on the core meaning, thereby aligning better with the objective of simplification.
Figure 4: The workflow of an iteration of PGO. We present the template of PGO in Section  B.2.
Perturbation Summarization Simplification Classification
C1
C2
C3
W1
W2
W3
S1
S2
S3
Table 2: Sensitivity of each task on different perturbations, ✓ for sensitive and ✗ for robust.
Specifically, we present the robustness relationship between each task and each perturbation. The detailed results are shown in the Table 2. In the subsequent construction of robust prompts, we focus only on perturbations that lead to performance degradation in LLMs.
3 Prompt Generation Framework
To enhance auto-prompt robustness under perturbations, we propose PGO, a real gradient-free optimization framework. Unlike traditional methods requiring gradient access, PGO leverages pseudo guidance signals to iteratively generate and optimize prompts, improving robustness in black-box LLMs settings.
3.1 Analyzing General Method
To enhance prompt robustness for task performance, many researchers have turned to data augmentation, adding perturbed text to the training data and training the model on this augmented dataset to produce robust prompts. We conduct an ablation study on dataset with perturbed data.The details are in Experiments. The result shows, in some tasks, the performance of these prompts even declines, likely due to the excessive diversity of perturbation types introduced during augmentation. We contend that while basic data augmentation methods can be beneficial in certain contexts, they may inadvertently introduce noise that interferes with effective prompt generation and limits generalization to unseen perturbations.
3.2 Framework of PGO
PGO draws on the idea of adversarial training. We divide the process of optimization in two phases and uses a pseudo gradient to substitute the real gradient to add perturbation and optimization. It first iteratively adds perturbations within the specified range to undisturbed inputs by adjusting them along the gradient direction to maximize disruption. The perturbed samples generated are then used for training based on a defined loss function. The workflow of our method is shown in Figure 4. PGO contains many iterations and every iterations include two key components:
  • Pseudo Gradient Perturbation Phase. PGO begins with manually crafted prompts and unperturbed text as the initial input. In Section.2, we categorize nine distinct perturbations into two types and design tailored perturbation methods for each category. These methods generate perturbed samples that simulate various perturbation scenarios, allowing the subsequent optimization algorithms to build robustness across all types of perturbations. This approach ensures that the model remains effective when faced with diverse perturbations.
  • Pseudo Gradient Optimization Phase. At this stage, we employ an iterative optimization method that incorporates a "gradient" mechanism to guide the LLMs in refining the prompts. For each generated prompts, we select the candidate with the best performance on the validation set and retain it for the next iteration.
3.3 Pseudo Gradient Perturbation Phase